# Protap: A Benchmark for Protein Modeling on Realistic Downstream Applications

### v2

@@ 4.2 Impacts of Pretraining Strategies and Architectures on Downstream Performance
Pear(%$\uparrow$ 57.11 $\pm$ 1.11 55.91 $\pm$ 1.24 51.15 $\pm$ 2.00 55.82 $\pm$ 1.67 60.92 $\pm$ 0.22 70.97 $\pm$ 1.82 50.30 $\pm$ 3.60
Table 8: Comparisons between general and domain models on protein function prediction.
General Architecture Domain Model
Metrics EGNN w/PFP SE3 GVP w/PFP ProtBERT w/PFP D-Transformer DeepFRI Deepfunc
Fmax(%$\uparrow$ 9.06 $\pm$ 0.18 3.93 $\pm$ 0.00 10.12 $\pm$ 0.07 6.05 $\pm$ 0.05 19.57 $\pm$ 0.80 26.01 $\pm$ 0.45 45.40 $\pm$ 0.41
AUPR(%$\uparrow$ 8.23 $\pm$ 0.09 4.10 $\pm$ 0.23 5.30 $\pm$ 0.02 5.64 $\pm$ 0.01 11.16 $\pm$ 0.67 32.77 $\pm$ 0.58 50.84 $\pm$ 0.30
4.3 Comparison Between Pretrained Models and Domain-Specific Models
@@
We compare the pretraining model architectures with eight domain-specific models, which are listed in Tab. 2. Since the representative solution of mutation effect prediction is protein language models, we omit the comparison on mutation effect prediction. For the pretraining model architectures, we report the best training strategies according to Tab. 4. The comparison between pretrained model architectures and domain-specific architectures is given in Tab. LABEL:tab:enzyme_domain-8. In particular, we aim to answer the following research questions.
@@
RQ4: How do the domain-specific models perform compared with general ones?
@@
The performance of domain-specific models and pretrained models varies across different tasks. We observe that on the protein-ligand interaction and protein function annotation prediction task, the domain-specific models could outperform the general model architectures by a large margin. However, for the applications of PROTACs that exhibit complex interaction processes, the general framework EGNN exhibits great results.
@@
RQ5: To what extent does incorporating biochemical inductive biases improve generalization across Enzyme-Catalyzed Protein Cleavage Site Prediction?
EGNN, UniZyme, and D-Trans exhibit superior performance on the protein task, e.g., D-Trans achieves $97.60\pm 0.08$ on C14.005 while EGNN achieves $90.24\pm 0.32$ on M10.003. UniZyme obtains comparable results. These indicate that the use of biochemical priors, such as energy frustration and distance matrix, facilitates improved detection of functional regions in proteins.
5 Conclusion and Future Works
@@
We introduce Protap, a unified benchmark that brings together general-purpose pretraining architectures and domain-specific models to evaluate five key protein modeling applications under a standard, reproducible framework. Our extensive experiments indicate that no single model or pretraining objective can outperform all others across applications. The architecture and training strategy must be chosen based on the specific characteristics of each task. There are two directions that need further investigation: (i) We will further explore the scaling laws that govern how model capacity and pretraining data volume translate into downstream gains. (ii) We will extend Protap to holistically cover protein design tasks, e.g., peptide design, enzyme design.
Appendix A Supplemental Experiments and Observations
A.1 Descriptions and Categorization of Domain-Specific Models

### v3

@@ 5.1.2. Downstream Task Training Setup
EGNN From scratch 95.51 $\pm$ 0.03 88.55 $\pm$ 1.78 0.492 $\pm$ 0.031
Frozen encoder 95.11 $\pm$ 0.31 88.19 $\pm$ 1.59 0.482 $\pm$ 0.022
Multi-stage 95.72 $\pm$ 0.14 88.96 $\pm$ 1.42 0.501 $\pm$ 0.025
D-Trans From scratch 97.60 $\pm$ 0.08 86.58 $\pm$ 0.28 0.416 $\pm$ 1.028
Frozen encoder 96.53 $\pm$ 0.52 85.16 $\pm$ 0.89 0.445 $\pm$ 1.064
Multi-stage 98.11 $\pm$ 0.00 85.53 $\pm$ 0.61 0.425 $\pm$ 0.029
ProtBERT From scratch 95.49 $\pm$ 0.09 84.13 $\pm$ 1.89 0.521 $\pm$ 0.020
Frozen encoder 95.78 $\pm$ 0.21 87.46 $\pm$ 1.71 0.519 $\pm$ 0.033
Multi-stage 95.80 $\pm$ 0.02 88.00 $\pm$ 0.97 0.519 $\pm$ 0.017
ESM-2 Frozen encoder 97.23 $\pm$ 0.06 85.74 $\pm$ 3.47 0.491 $\pm$ 4.823
Multi-stage 98.16 $\pm$ 0.03 89.96 $\pm$ 1.49 0.404 $\pm$ 0.010
5.2. Analysis on Protein Model Pretraining
In this section, we examine three pretraining tasks employed in Protap as described in Sec. 3.2, which are MLM, MVCL, and PFP. We pre-train EGNN, SE(3) Transformer (SE(3) Trans), GVP, ProteiBERT (ProtBERT), and D-Transformer (D-Trans) with all three tasks with 542k proteins to obtain pretraining encoders, which are listed in Tab. 1 and are detailed in Appendix G. For ESM-2 and ESM-C, we testify to the publicly released pretrained weights that are learned via MLM training on 70M proteins. In particular, we aim to make comparisons across the following aspects:
RQ1 Scaling Law on Protein Modeling: Can large-scale pretraining encoders outperform a supervised encoder that trained on the small downstream training set from scratch?
Compared with Frozen Pretraining Encoders. The results of frozen pretraining encoders and supervised encoders are given in Table. 2, we observe that: (i) the unsupervised pretraining encoders can achieve competitive performance on various downstream tasks. For example, using frozen pretraining encoders of ESM-2 and SaProt often outperform supervised encoders (e.g., EGNN, SE(3) Trans, and ProtBert ) on downstream tasks. (ii) However, current supervised encoders trained from scratch tend to outperform the unsupervised pretraining encoders even when they are trained with large-scale datasets. For example, the supervised encoder of D-Trans generally outperforms all other unsupervised encoders. This implies a degree of mismatch between the pretraining objectives and the downstream tasks, indicating that training from scratch enables the model to learn task-specific representations that are more aligned with the requirements of the downstream task.
Compared with Finetuned Pretraining Encoders. As shown in Tab. 3, by including the results of multi-stage fine-tuning in comparison, we find that multi-stage fine-tuning almost consistently outperforms the from-scratch and frozen-encoder setting, and the benefit becomes more pronounced as interaction complexity increases, particularly in PCS and PROTACs, where enzyme-substrate contact and ternary alignment require coordinating multiple interacting entities. These results indicate that for protein modeling tasks involving cross-modal alignment, effective use of pretrained models requires not only preserving pretrained knowledge but also providing the necessary representation adaptation capacity.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
(a) Parameter scaling
(b) PLI task data scaling
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
Figure 5. Scaling behavior in parameters and training data
Analysis of Scaling Law on Proteins. We analyze protein scaling laws from three perspectives. (i) Pre-train data scaling. As shown in Fig. 5(a), increasing the scale of pre-training data provides limited gains for specialized tasks. Models pre-trained on billions of sequences, such as the ESM family, do not consistently outperform those trained on much smaller corpora, and may even exhibit degraded performance. In contrast, large-scale pre-training yields clear improvements on general tasks, likely reflecting enhanced capture of protein evolutionary signals. (ii) Model parameter scaling. Fig. 5(a) reveals a pronounced model-size scaling law for general tasks, with performance steadily improving as model size increases from 2M to 650M parameters (e.g., ESM2 vs. ProteinBERT). However, this trend largely vanishes on specialized tasks such as PROTACs. (iii) Domain data scaling. Fig. 5(b) shows that downstream data scale also exhibits a power-law behavior: expanding the PLI dataset from approximately 14k to nearly 100k samples leads to substantial performance gains, which highlights that task-specific data plays a vital role in realizing scaling benefits.
RQ2 Sequential or Structural Models: Are models incorporating structures better than those purely using protein sequences?
As shown in Fig. 6, we observe that on specialized applications, the models with lower average ranks are all structure-aware models. For general tasks such as PFA and MTP that rely more on evolutionary and semantic statistical regularities, large-scale PLMs trained on much larger corpora show clear advantages. However, when controlling for training data scale and model capacity (E.g., ProtBERT vs. others), most structure-aware models still outperform the sequence-only model. These observations highlight two key insights: (i) incorporating three-dimensional structural information introduces essential inductive biases that are missing in sequence-only architectures. (ii) For tasks where performance depends strongly on geometric constraints, binding interfaces and pockets, and complex interactions, structural information can be more important than parameter scale to some extent.
RQ3 Pretraining Tasks: Is there any pretraining task that consistently outperforms others on downstream tasks?
To quantitatively assess whether any pretraining objective consistently outperforms others, we conduct rank-based Friedman tests with Nemenyi post-hoc analyses across all five model architectures, comparing three pretraining strategies. The results show that although the overall Friedman test reaches significance ($\chi^{2}{=}6.10$, p=0.047p{=}0.047), the concordance is low (W=0.153W{=}0.153), indicating limited agreement across model–task blocks; accordingly, no pairwise difference is significant under Nemenyi (CD=0.741{=}0.741, $\alpha{=}0.05$). Per-task Friedman tests show no significant differences on PCS, PROTACs, or PLI. The sole exception is PFA, where PFP achieves the best average rank (Friedman $\chi^{2}{=}6.40$, p=0.041p{=}0.041) and significantly outperforms MVCL in the Nemenyi test ($|\Delta R|{=}1.60>\text{CD}{=}1.48$). Detailed results are presented in Tab. 9. Overall, these results suggest no universal advantage of a single pretraining objective, while PFP specifically benefits functional annotation.
A short description of the image A long description of the image content, detailing all important visual information for a reader using screen-reading software.
(a) Specialized applications
(b) General applications
Figure 6. Performance ranking of structure-aware and sequence-only models on downstream applications.
5.3. Analysis on Domain Models
@@
We compare the pretraining model architectures with eight domain-specific models, which are listed in Tab. 1. Since the representative solution of mutation effect prediction is protein language models, we omit the comparison on mutation effect prediction. For the pretraining model architectures, we report the best-performing training strategies according to Tab. 2. The comparison between pretrained model architectures and domain-specific architectures is presented in Fig. 4 (a)–(d). In particular, we aim to answer the following research questions.
@@
RQ4: Domain Model or General Encoder: How do the domain-specific models perform compared with general ones?
@@
The performance of domain-specific models and pretrained models varies across different tasks. We observe that on the protein-ligand interaction and protein function annotation prediction task, the domain-specific models could outperform the general model architectures by a large margin. However, for the applications of PROTACs that exhibit complex interaction processes, the general framework EGNN exhibits great results. SaProt greatly outperforms other models in mutation effects prediction, including ESM-2. This is likely attributable to the Structure-Vocabulary, which reduces local sequence ambiguity and mitigates semantic confusion.
@@
RQ5: Effect of Biochemical and Structural Priors: To what extent does incorporating biochemical or structural inductive biases improve generalization across downstream applications?
We observe that several domain-specific models outperform, or are comparable to, general-purpose models, including UniZyme, DeepFunc, DeepFRI, and KDBNet. In contrast, models relying only on 2D cues or sequence features, such as MONN, DeepPROTACs, and ET-PROTACs, consistently perform the worst. Notably, the strong domain models share two common traits: they leverage pretrained representations for initialization or feature enhancement, and they incorporate task-aligned priors, such as family information for PFA and energy-frustration signals for PCS. These findings suggest that pretrained representations can effectively strengthen domain models, while prior knowledge should be carefully chosen to match the data and task.
6. Conclusion and Future Works
@@
We introduce Protap, a unified benchmark that brings together general-purpose pretraining architectures and domain-specific models to evaluate five key protein modeling applications under a standard, reproducible framework. Our extensive experiments indicate that no single model or pretraining objective can outperform all others across applications. There are two directions that need further investigation: (i) We will further explore the scaling laws that govern how model capacity and pretraining data volume translate into downstream gains. (ii) We will extend Protap to holistically cover protein design tasks, e.g., peptide design, enzyme design.
Appendix A Abbreviation Statement
The Tab.4 lists the abbreviations used throughout the paper, covering downstream applications, pretraining objectives, model architectures, datasets, and biological terminology.
Table 4. List of abbreviations used in this paper.
Abbreviation Description
Downstream Applications
PCS Protein Cleavage Site prediction
PROTACs Proteolysis-Targeting Chimeras
PLI Protein–Ligand Interaction
PFA Protein Function Annotation
MTP Mutation Effect Prediction
Pretraining Objectives
MLM Masked Language Modeling
MVCL Multi-View Contrastive Learning
PFP Protein Family Prediction
Models & Architectures
EGNN Equivariant Graph Neural Network
SE(3) Trans SE(3)-Equivariant Transformer
GVP Geometric Vector Perceptron
D-Trans Distance-aware Transformer
ProtBERT ProteinBERT
ESM-C ESM Cambrian
Datasets & Resources
AFDB AlphaFold Protein Structure Database
UR50 UniRef50 Protein Sequence Cluster
**Biological terminology**
GO Gene Ontology
Appendix B Reproducibility
Our implementations are based on widely adopted architectures, pretraining strategies, and downstream evaluation pipelines for protein modeling. The code is available at https://github.com/Trust-App-AI-Lab/protap. To ensure reproducibility, we provide detailed descriptions of the pretraining objectives in Section 3.2, along with the corresponding training configurations in Appendix G.2. Hyperparameters for all models, including learning rate schedules, batch sizes, and optimization settings, are explicitly reported in Section 5 and Appendix G.2. The datasets used for both pretraining and downstream evaluation are publicly available at https://huggingface.co/datasets/findshuo/Protap, and we specify their sources in Tab. 1 (b), and provide more details in Appendix F. The exact evaluation metrics are specified in Section Appendix F. These details, combined with the systematic comparison across architectures and strategies presented in Tab. 2 and Fig. 4, allow independent researchers to reproduce our results based on the manuscript alone.
Appendix C Comparison of benchmark coverage for protein modeling
We compare across three dimensions: applications (specialized vs. general), pretraining tasks, and model architectures (domain-specific vs. general). Comparisons with ProteinGym and ProteinBench are not included because their focuses differ from Protap. ProteinGym (Notin et al., 2023) primarily targets mutation effect prediction with task-specific models. ProteinBench (YE et al., 2025) emphasizes generative design tasks, e.g., backbone design and sequence-structure co-design, which are currently beyond Protap’s scope. Detailed objective coverage of benchmarks is provided in Tab.5.
Table 5. Comparison of benchmark coverage for protein modeling. indicates the presence of this dimension, while ○ denotes its absence.
Protap (Ours) PEER (Xu et al., 2022) ProteinWorkshop (Jamasb et al., 2024) TAPE (Rao et al., 2019) ProteinGLUE (Capel et al., 2022)
Targeted protein degradation by Proteolysis-targeting chimeras
Protein–ligand interactions
General Applications Function prediction
Mutation effect prediction
Pretraining Tasks Masked language modeling
Multi-view contrastive learning
Protein family prediction
General Architectures Protein language models
Geometric GNNs
Sequence-structure hybrid models
Domain-specific Architectures Protein function models
Enzyme domain models
Ternary complexes models
Appendix D Supplemental Experiments and Observations
Table 6. Performance comparison across model architectures under different training strategies with full parameter fine-tuning.
Model PCS PROTACs PLI
C14.005 PROTACDB Davis
[…]
EGNN w/MLM 95.36 $\pm$ 0.06 19.23 $\pm$ 0.31 79.03 $\pm$ 0.61 88.50 $\pm$ 0.37 0.461 $\pm$ 0.010 57.56 $\pm$ 1.18
EGNN w/MVCL 94.80 $\pm$ 0.00 19.61 $\pm$ 1.16 76.89 $\pm$ 2.50 88.81 $\pm$ 0.50 0.484 $\pm$ 0.031 56.18 $\pm$ 0.12
EGNN w/PFP 95.09 $\pm$ 0.41 27.95 $\pm$ 0.99 79.01 $\pm$ 0.55 87.65 $\pm$ 0.33 0.523 $\pm$ 0.036 55.44 $\pm$ 2.72
SE(3) Trans w/MLM 67.65 $\pm$ 1.31 0.55 $\pm$ 0.03 77.54 $\pm$ 0.06 87.08 $\pm$ 1.79 0.681 $\pm$ 0.109 43.20 $\pm$ 4.63
SE(3) Trans w/MVCL 44.01 $\pm$ 0.72 0.20 $\pm$ 0.00 75.59 $\pm$ 2.54 83.70 $\pm$ 0.04 1.110 $\pm$ 0.074 34.34 $\pm$ 4.76
SE(3) Trans w/PFP 69.27 $\pm$ 2.55 0.81 $\pm$ 0.17 77.60 $\pm$ 0.39 87.06 $\pm$ 0.66 0.536 $\pm$ 0.047 47.94 $\pm$ 3.51
ProtBERT w/MLM 95.45 $\pm$ 0.10 9.37 $\pm$ 0.62 77.09 $\pm$ 0.36 85.83 $\pm$ 0.73 0.511 $\pm$ 0.011 49.81 $\pm$ 1.32
ProtBERT w/MVCL 95.78 $\pm$ 0.21 28.51 $\pm$ 0.06 73.58 $\pm$ 0.13 83.40 $\pm$ 0.33 0.511 $\pm$ 0.008 50.25 $\pm$ 1.76
ProtBERT w/PFP 64.25 $\pm$ 0.35 0.36 $\pm$ 0.00 78.79 $\pm$ 0.85 87.41 $\pm$ 1.00 0.529 $\pm$ 0.018 49.99 $\pm$ 1.77
GVP w/MLM 89.09 $\pm$ 0.00 11.76 $\pm$ 0.01 72.54 $\pm$ 0.06 79.03 $\pm$ 0.01 0.438 $\pm$ 0.013 47.28 $\pm$ 0.00
GVP w/MVCL 88.68 $\pm$ 0.00 10.96 $\pm$ 0.01 71.45 $\pm$ 0.42 78.19 $\pm$ 0.13 0.423 $\pm$ 0.011 59.31 $\pm$ 0.00
GVP w/PFP 71.83 $\pm$ 0.01 1.50 $\pm$ 0.00 72.72 $\pm$ 0.36 79.94 $\pm$ 0.58 0.431 $\pm$ 0.002 59.04 $\pm$ 0.00
D-Trans w/MLM 94.40 $\pm$ 1.13 29.93 $\pm$ 0.09 74.46 $\pm$ 0.78 81.44 $\pm$ 0.59 0.464 $\pm$ 0.023 55.79 $\pm$ 0.50
D-Trans w/MVCL 98.11 $\pm$ 0.28 36.50 $\pm$ 3.17 73.26 $\pm$ 0.18 81.25 $\pm$ 0.50 0.437 $\pm$ 0.002 59.91 $\pm$ 0.13
D-Trans w/PFP 96.65 $\pm$ 0.21 34.55 $\pm$ 2.53 73.50 $\pm$ 0.30 82.64 $\pm$ 0.05 0.441 $\pm$ 0.020 59.66 $\pm$ 0.75
ESM-2 93.10 $\pm$ 1.71 55.09 $\pm$ 0.22 81.25 $\pm$ 1.44 88.17 $\pm$ 1.10 0.416 $\pm$ 0.011 59.59 $\pm$ 0.35
[…]
SaProt 94.75 $\pm$ 0.00 57.35 $\pm$ 0.01 81.43 $\pm$ 0.90 87.16 $\pm$ 0.13 0.423 $\pm$ 0.008 58.81 $\pm$ 0.02
Table 7. Performance comparison across model architectures under different fine-tuning strategies.
Model Pretrain objective LR PCS PROTACs PLI
C14.005 PROTACDB Davis
[…]
EGNN MLM 5e-6 $95.49\pm 0.44$ $32.61\pm 1.09$ $78.98\pm 2.09$ $87.02\pm 1.53$ $0.497\pm 0.042$ $52.80\pm 1.84$
1e-6 $95.47\pm 0.23$ $34.34\pm 0.70$ $0.01\pm 1.92 88$$8.96\pm 1.24$ $0.485\pm 0.018$ $56.50\pm 3.65$
1e-7 $95.41\pm 0.37$ $28.14\pm 1.19$ $79.11\pm 1.69$ $87.77\pm 1.16$ $0.506\pm 0.051$ $52.92\pm 3.82$
MVCL 5e-6 $95.24\pm 0.36$ $32.91\pm 0.08$ $79.33\pm 2.59$ $86.81\pm 0.90$ $0.491\pm 0.027$ $52.91\pm 5.00$
1e-6 $95.72\pm 0.14$ $35.04\pm 1.11$ $79.47\pm 2.25$ $86.88\pm 1.07$ $0.503\pm 0.039$ $52.86\pm 2.35$
1e-7 $95.17\pm 0.04$ $30.05\pm 1.04$ $79.70\pm 2.36$ $87.79\pm 1.02$ $0.482\pm 0.022$ $52.77\pm 3.27$
PFP 5e-6 $95.01\pm 0.09$ $32.88\pm 0.42$ $79.52\pm 2.46$ $87.94\pm 1.46$ $0.508\pm 0.045$ $55.36\pm 3.59$
1e-6 $94.88\pm 0.39$ $28.63\pm 1.06$ $79.88\pm 1.80$ $88.36\pm 1.39$ $0.494\pm 0.033$ $56.39\pm 3.48$
1e-7 $94.87\pm 0.34$ $32.69\pm 0.10$ $78.87\pm 2.51$ $87.13\pm 1.02$ $0.499\pm 0.015$ $53.31\pm 2.46$
D-Trans MLM 5e-6 98.11 $\pm$ 0.00 63.50 $\pm$ 0.18 77.76 $\pm$ 2.55 85.15 $\pm$ 1.42 0.450 $\pm$ 0.040 59.59 $\pm$ 0.91
1e-6 97.88 $\pm$ 0.07 60.12 $\pm$ 0.32 76.08 $\pm$ 1.19 83.46 $\pm$ 0.64 0.465 $\pm$ 0.034 58.07 $\pm$ 0.53
1e-7 97.76 $\pm$ 0.09 57.79 $\pm$ 0.87 74.88 $\pm$ 0.85 82.46 $\pm$ 0.82 0.467 $\pm$ 0.022 56.71 $\pm$ 0.14
MVCL 5e-6 96.99 $\pm$ 0.06 72.43 $\pm$ 0.47 72.12 $\pm$ 4.08 84.09 $\pm$ 0.88 0.670 $\pm$ 0.045 57.31 $\pm$ 0.53
1e-6 97.27 $\pm$ 0.07 72.56 $\pm$ 0.04 74.70 $\pm$ 1.44 82.29 $\pm$ 0.21 0.549 $\pm$ 0.032 55.51 $\pm$ 0.29
1e-7 97.32 $\pm$ 0.07 71.68 $\pm$ 0.04 74.22 $\pm$ 1.44 82.02 $\pm$ 0.25 0.514 $\pm$ 0.026 64.59 $\pm$ 0.64
PFP 5e-6 96.93 $\pm$ 0.05 67.96 $\pm$ 0.25 77.46 $\pm$ 0.76 85.53 $\pm$ 0.61 0.425 $\pm$ 0.029 58.95 $\pm$ 2.09
1e-6 97.17 $\pm$ 0.03 65.49 $\pm$ 0.64 75.00 $\pm$ 0.51 83.69 $\pm$ 0.51 0.435 $\pm$ 0.031 57.41 $\pm$ 2.27
1e-7 97.29 $\pm$ 0.01 63.70 $\pm$ 1.21 74.40 $\pm$ 0.51 83.28 $\pm$ 0.47 0.443 $\pm$ 0.028 56.36 $\pm$ 2.10
ProtBERT MLM 5e-6 $95.80\pm 0.02$ $32.37\pm 0.11$ $76.95\pm 2.35$ $84.53\pm 1.31$ $0.530\pm 0.017$ $54.84\pm 1.84$
1e-6 $95.50\pm 0.14$ $22.63\pm 0.48$ $80.21\pm 2.41$ $86.84\pm 1.39$ $0.519\pm 0.034$ $53.99\pm 1.21$
1e-7 $95.46\pm 0.24$ $15.27\pm 0.56$ $79.72\pm 1.92$ $85.10\pm 1.37$ $0.554\pm 0.021$ $53.04\pm 1.75$
MVCL 5e-6 $95.74\pm 0.21$ $23.83\pm 0.14$ $78.05\pm 1.21$ $84.83\pm 1.88$ $0.546\pm 0.035$ $55.28\pm 1.91$
1e-6 $95.25\pm 0.20$ $33.23\pm 0.53$ $80.42\pm 1.59$ $85.53\pm 1.45$ $0.534\pm 0.026$ $55.73\pm 1.29$
1e-7 $95.05\pm 0.09$ $33.25\pm 0.65$ $79.43\pm 1.55$ $86.98\pm 1.68$ $0.547\pm 0.035$ $55.78\pm 1.63$
PFP 5e-6 $64.52\pm 0.44$ $4.54\pm 0.80$ $78.35\pm 1.89$ $84.69\pm 1.39$ $0.527\pm 0.034$ $52.34\pm 1.62$
1e-6 $63.88\pm 0.25$ $6.41\pm 0.60$ $77.85\pm 1.82$ $88.00\pm 0.97$ $0.553\pm 0.033$ $53.82\pm 1.77$
1e-7 $64.24\pm 0.38$ $0.41\pm 0.60$ $78.94\pm 2.72$ $86.56\pm 1.35$ $0.549\pm 0.037$ $51.36\pm 1.29$
ESM-2 MLM 5e-6 94.77 $\pm$ 0.12 28.30 $\pm$ 0.44 82.09 $\pm$ 2.21 89.96 $\pm$ 1.49 0.404 $\pm$ 0.010 61.04 $\pm$ 0.03
1e-6 96.25 $\pm$ 0.11 32.35 $\pm$ 0.37 81.37 $\pm$ 1.19 89.31 $\pm$ 1.60 0.404 $\pm$ 0.010 61.03 $\pm$ 0.04
1e-7 98.16 $\pm$ 0.03 38.14 $\pm$ 0.38 81.25 $\pm$ 1.19 88.99 $\pm$ 1.59 0.405 $\pm$ 0.011 61.01 $\pm$ 0.02
D.1. Descriptions and Categorization of Domain-Specific Models